OpenStudy (anonymous):

what is y=-7x+4 and x=-2 by substitution

OpenStudy (anonymous):

y=-7(-2)+4 y=-10

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

wats 5>9+r and how would it look on a number line

OpenStudy (gw2011):

The answer is y=18. y=(-7)(-2)+4 y=14+4 y=18

OpenStudy (anonymous):

is that the answer to the first or second question

OpenStudy (anonymous):

y=18 is the correct answer, not y=10. when you multiply 2 negatives, it will become a positive, making it +14+4=18. For your second question, 5>9+r. Simply subtract 9 from both sides making it -4>r. Then to put it on a number line, just put an open circle on x=-4 and indicate that r is everything that is smaller than than -4, so shade the number line to the left for all negative numbers less than -4.

OpenStudy (anonymous):

thank you i have another qustion

OpenStudy (anonymous):

can you please help me

OpenStudy (anonymous):

sorry, yea sure